<p>15:30</p>
<p>1.5 MeV/u beam currently being tuned</p>
<p>Target is ON, setpoint = 135 K</p>
<p>System checks OK - white rabbit fail diff ~ 9</p>
<p>Temperatures OK</p>
<p>Merger OK</p>
<p>Data rate ~ 1400 kb/sec</p>
<p>Leakage current ~ 3.41 uA</p>
<p>Aida 1 and 2 missing channels from previous days appear working on stat layout 2 and xy pixel total plot (attachment 3)</p>
<p>17:12</p>
<p>Beam tuning complete. See attached Ex rate histogram.</p>
<p>The cryo compressor is being turned OFF for around 1 hour to remove any frozen nitrogen which may be the cause of the increased pressures observed.</p>
<p>Beam OFF to do so.</p>
<p>System checks OK - white rabbit fail diff ~ 9</p>
<p>Temperatures OK</p>
<p>Merger OK</p>
<p>Data rate ~ 200 kb/sec</p>
<p>Compressor OFF, setpoint on controller is 273 K</p>
<p>18:51</p>
<p>Pressure in the ring recovered quickly after closing valves to gas jet (attachment 5)</p>
<p>Increasing temperature caused a large spike in pressure in the inlets. E1 turbo briefly shut down due to the pressure increase, since restarted and pressure back in the -4 mbar range. (attachment 6)</p>
<p>Temperature being decreased</p>
<p>Beam back ON</p>
<p>data rate ~ 1500 kb/sec</p>
<p>merger OK</p>
<p>Temperatures OK</p>
<p>System checks OK - white rabbit diff ~ 12</p>
<p>21:58</p>
<p>After cooling the jet target, expected density is not achieved. E4 pressure is ~4E-10 mbar compared to ~2E-8mbar previously. Temperature reduced to 127K E4 pressure increases however stilll within -10 mbar range.</p>
<p>With beam on target we see rutherford peak at ~3 MeV, however it is much reduced.</p>
<p>Beam OFF to investigate inside the cave.</p>
<p>Decreasing temperature of nozzle in 0.5 k steps to see effect on E4 and S1 dump.</p>
<p> </p> |